OpenStudy (anonymous):

a ramp with a 5.4 metre horizontal length rising at 5 degrees what is the height of the ramp at its highest point

OpenStudy (anonymous):

let x be height of the highest point then sin 5 =x/5.4 x= sin 5 * 5.4 m =0.47m